OpenStudy (anonymous):

In 1997 there were 14 large apartment buildings in Calston, TX. By 2010 there were 53 large apartment buildings. What is the rate of change in the number of apartments buildings from 1997 to 2010? A. 1/39 years per building B. 53/13 buildings per year C. 3 buildings per year D. 14 buildings per month

OpenStudy (anonymous):

2010-1997=13 years 53-14 = 39 39/13 = 3 3 buildings per year

OpenStudy (agent0smith):

^yep Rate of change = change in y divided by change in x note that time will ~always be the x.

OpenStudy (anonymous):

thank you to both of you:)